luis begins walking at speed of 4mph, 45mins later his wife notices he left his wallet so gets on her bike and begins riding at 24mph along the same path luis took
How long will it take for her to catch luis?
How far from their house will she catch luis?

Answer by lwsshak3(11628)

luis begins walking at speed of 4mph, 45mins later his wife notices he left his wallet so gets on her bike and begins riding at 24mph along the same path luis took
How long will it take for her to catch luis?
How far from their house will she catch luis?
**
45min=.75 hr
let x=minutes wife will travel to catch luis
4(.75+x)=distance wife will travel to catch luis
distance/speed=travel time
4(.75+x)/24=x
24x=3+4x
20x=3
x=3/20 hr=.15 hr=9 min
4(.75+x)=4(.75+.15)=4*.9=3.6 mi
ans:
It will take wife 9 minutes to catch luis.
She will be 3.6 mi from the house when she catches up with luis